OpenStudy (anonymous):

can someone show me the steps to the answer, i have the answer but need to see it written out. 4/x=9/5-7x-4/5x

OpenStudy (anonymous):

\[\frac{ 4 }{ x }=\frac{ 9 }{ 5 }-\frac{ 7x-4 }{ 5x }\]

OpenStudy (anonymous):

answer of 8